# Trial and Trust: Addressing Byzantine Attacks with Comprehensive Defense Strategy

### v1

@@ Appendix G Local methods
$+3L^{2}\gamma^{3}(1+3L\gamma)\sigma^{2}(l-1)+\frac{3L\gamma^{2}}{2G}\sigma^{2}+\gamma(2+3L\gamma)\delta_{1}+\gamma\zeta(N)+\delta.$
@@
$E\left[\hat{f}\left(x^{T}\right)-\hat{f}\left(x^{0}\right)\right]$$\leqslant$ $-\frac{\gamma}{2}\left(\left(1-4\delta_{2}\right)-3L\gamma(1+2\delta_{2})\right)\sum\limits_{t=0}^{T-1}\mathbb{E}\left\|\nabla f\left(\overline{x}^{t}\right)\right\|^{2}$
$+\frac{9\delta_{2}L\gamma^{2}(1+3L\gamma)(l-1)}{2}\sum\limits_{t=0}^{T-1}\mathbb{E}\left\|\nabla f(\overline{x}^{t})\right\|^{2}+\frac{9\delta_{1}L\gamma^{2}(1+3L\gamma)(l-1)T}{2}$
$+3L^{2}\gamma^{3}(1+3L\gamma)\sigma^{2}(l-1)T+\frac{3L\gamma^{2}T}{2G}\sigma^{2}+\gamma(2+3L\gamma)\delta_{1}T+\gamma\zeta(N)T+\delta T$
@@
$+3L^{2}\gamma^{3}(1+3L\gamma)\sigma^{2}(l-1)T+\frac{3L\gamma^{2}T}{2G}\sigma^{2}+\frac{9\delta_{1}L\gamma^{2}(1+3L\gamma)(l-1)T}{2}$
$+\gamma(2+3L\gamma)\delta_{1}T+\gamma\zeta(N)T+\delta T.$
@@
$1-4\delta_{2}-3L\gamma(1+2\delta_{2})-9L\gamma\delta_{2}(1+3L\gamma)(l-1)$$\geqslant$ $\frac{1}{2}-\frac{15}{4}L\gamma-\frac{9}{8}L\gamma\left(1+\frac{3}{25(l-1)}\right)(l-1)$
@@
$+\frac{5}{2}\gamma\delta_{1}T+6L\gamma^{2}\delta_{1}(l-1)T+\gamma\zeta(N)+\delta T,$
@@
$+13\delta_{1}+30L\gamma(l-1)\delta_{1}+5\zeta(N)+5\frac{\delta}{\gamma},$
@@
In this remark we want to explain why $\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}\left\|\nabla f\left(\frac{1}{G}\sum\limits_{i\in\mathcal{G}}x_{i}^{t}\right)\right\|^{2}$, that we choose as a criterion in Theorem G.3 is consistent. When we consider iterations, when Algorithm 7 performs aggregation, we assign weights of honest devices, i.e. devices, which send honest stochastic gradient $g$ with $\mathbb{E}[g]=\nabla f(x)$ equal to $\frac{1}{G}$, and weights of other devices (who acts as Byzantine) is equal to 00. And this layout gives an estimate that is an upper bound of true $\omega$ realization (G.2). In other words, we say, that if algorithm at each aggregation round take average of points from devices, which all previous local round act like honest it would not be better than for convergence, than real iteration of the algorithm. In such a way, we a not interesting in the points of devices, who perform even one Byzantine-like iteration in the local round. Now it is clear why we have to weaken the assumption about at least one honest device at each iteration, not necessary the same: we request at least one honest device at each local round, not necessary the same in different rounds.
@@
[…]
@@ Appendix H Partial participation
In the previous section, we discussed an important regime in distributed systems known as the local approach. Another widely used scenario in distributed learning is partial participation, which can be advantageous in various practical setups (Yang et al., 2021; Wang & Ji, 2022; Li et al., 2022). In this section, we extend our AutoBANT algorithm to support partial participation in federated learning. Below, we provide a formal description of the AutoBANT with Partial Participation method (Algorithm 8).
@@
Under Assumptions 2.1, 2.2(c), 2.3, 2.4 with $\delta_{2}\leqslant 0.25$, 2.5, for solving the problem (1), after $T$ iteration of Algorithm 2 in partial participation scenario with $\gamma\leqslant\frac{1}{15L}$, it implies
$\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}\|\nabla f(x^{t})\|^{2}$$\leqslant$ $\frac{4\mathbb{E}\left[\hat{f}\left(x^{0}\right)-\hat{f}\left(\hat{x}^{*}\right)\right]}{\gamma T}+3\delta_{1}+\frac{6L\gamma}{\widetilde{G}}\sigma^{2}+4\zeta(N)+\frac{4\delta}{\gamma},$
@@
$\hat{f}(x^{t+1})$$\leqslant$ $\min_{\omega\in\Delta_{1}^{n}}\hat{f}\left[x^{t}-\gamma\sum_{i\in\mathcal{W}(t)}\omega_{i}g_{i}^{t}\right]+\delta$
$\leqslant$ $\hat{f}\left[x^{t}-\frac{\gamma}{G(t)}\sum_{i\in\mathcal{G}(t)}g_{i}^{t}\right]+\delta$
(${\leqslant}$ $\hat{f}(x^{t})-\left\langle\nabla\hat{f}(x^{t}),\frac{\gamma}{G(t)}\sum_{i\in\mathcal{G}(t)}g_{i}^{t}\right\rangle+\frac{L\gamma^{2}}{2}\left\|\frac{1}{G(t)}\sum_{i\in\mathcal{G}(t)}g_{i}^{t}\right\|^{2}+\delta$
@@
$\mathbb{E}\hat{f}(x^{t+1})$$\leqslant$ $\mathbb{E}\hat{f}(x^{t})-\left\langle\nabla\hat{f}(x^{t}),\frac{\gamma}{G(t)}\sum_{i\in\mathcal{G}(t)}\nabla f_{i}(x^{t})\right\rangle+\frac{L\gamma^{2}}{2}\mathbb{E}\left\|\frac{1}{G(t)}\sum_{i\in\mathcal{G}(t)}g_{i}^{t}\right\|^{2}+\delta$
@@
$+\frac{3\gamma}{2}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}\right)+\frac{L\gamma^{2}}{2}\mathbb{E}\left\|\frac{1}{G(t)}\sum_{i\in\mathcal{G}(t)}g_{i}^{t}\right\|^{2}+\delta$
(${\leqslant}$ $\mathbb{E}\hat{f}(x^{t})+\gamma\zeta(N)-\frac{\gamma}{2}\left\|\nabla f(x)\right\|^{2}+\frac{3\gamma}{2}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}\right)$
$+\frac{3L\gamma^{2}}{2}\left\|\frac{1}{G(t)}\sum_{i\in\mathcal{G}(t)}(\nabla f(x^{t})-\nabla f_{i}(x^{t}))\right\|^{2}+\frac{3L\gamma^{2}}{2}\mathbb{E}\left\|\frac{1}{G(t)}\sum_{i\in\mathcal{G}(t)}(\nabla f_{i}(x^{t})-g_{i}^{t})\right\|^{2}$
$+\frac{3L\gamma^{2}}{2}\left\|\frac{1}{G(t)}\sum_{i\in\mathcal{G}(t)}\nabla f(x^{t})\right\|^{2}+\delta.$
@@
$\mathbb{E}\hat{f}(x^{t+1})$ (${\leqslant}$ $\mathbb{E}\hat{f}(x^{t})+\gamma\zeta(N)-\frac{\gamma}{2}\left\|\nabla f(x)\right\|^{2}+\frac{3\gamma}{2}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}\right)$
$+\frac{3L\gamma^{2}}{2}\left(\frac{1}{G(t)}\sum_{i\in\mathcal{G}(t)}\left\|\nabla f(x^{t})-f_{i}(x^{t})\right\|^{2}+\frac{1}{(G(t))^{2}}\sum_{i\in\mathcal{G}(t)}\mathbb{E}\left\|\nabla f_{i}(x^{t})-g_{i}^{t}\right\|^{2}\right)$
$+\frac{3L\gamma^{2}}{2}\left\|\nabla f(x^{t})\right\|^{2}+\delta$
@@
$+\frac{3L\gamma^{2}}{2}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}+\frac{\sigma^{2}}{G(t)}\right)+\frac{3L\gamma^{2}}{2}\left\|\nabla f(x^{t})\right\|^{2}+\delta$
@@
$+\frac{2\gamma}{2}(1+3L\gamma)\delta_{1}+\frac{3L\gamma^{2}}{2G(t)}\sigma^{2}+\gamma\zeta(N)+\delta.$
@@
$\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}\|\nabla f(x^{t})\|^{2}$$\leqslant$ $\frac{4\mathbb{E}\left[\hat{f}\left(x^{0}\right)-\hat{f}\left(\hat{x}^{*}\right)\right]}{\gamma T}+3\delta_{1}+\frac{6L\gamma}{\widetilde{G}}\sigma^{2}+4\zeta(N)+\frac{4\delta}{\gamma}.$
@@
[…]
@@ Appendix I SimBANT
We also propose another method for dealing with Byzantine attacks - SimBANT. Here we use the concept of the trial function in a different way: the key is that we look not at the loss function $\hat{f}$, as in Algorithms 1, 2, but at the output of the model $m$ on the trial data $\hat{D}$. In particular, the calculation of trust scores $w_{i}$ is now based on how the outputs of the model parameters obtained on the server (a guaranteed honest device, without losing generality, we can assume that it has index 1) and on the device are similar to each other. The server is validated in pairs with each device and the trust scores are calculated for each device relative to the prediction classes based on how similar they are between the server and the device. And only on the basis of the trust scores, we give a weights to devices’ trained models and calculate the new state of the general model. Let us denote by $g^{t}$ the model that the server sends to the devices at step $t$, $g_{i}^{t},i\in\{2,\ldots,n\}$ devices’ trained models, $g_{1}^{t}$ the trained server model. Then the trust score for $i$-th device is function $\alpha_{i}\rightarrow sim(m(x^{t}-\gamma g_{i}^{t},\mathcal{\hat{D}}),m(x^{t}-\gamma g_{1}^{t},\mathcal{\hat{D}}))$, which measures the similarity of predictions of server and device models, where $x^{t}-\gamma g_{i}^{t}$ and $x^{t}-\gamma g_{1}^{t}$ are new parameters on the $i$-th device and the server, respectively, $sim$ is responsible for the measure of the similarity of the outputs. We do not provide the theory for this method but validate it in practice (see Section 5).
@@
10:     $\omega_{i}^{t}=(1-\beta)\omega_{i}^{t-1}+\beta\frac{sim(m(x^{t}-\gamma g_{i}^{t},\hat{D}),m(x^{t}-\gamma g_{1}^{t},\hat{D}))}{\sum\limits_{j=1}^{n}sim(m(x^{t}-\gamma g_{j}^{t},\hat{D}),m(x^{t}-\gamma g_{1}^{t},\hat{D}))}$
11:     $x^{t+1}=x^{t}-\gamma\sum\limits_{i=1}^{n}\omega_{i}^{t}g_{i}^{t}$

### v5

@@ Appendix G Local methods
$+3L^{2}\gamma^{3}(1+3L\gamma)\sigma^{2}(l-1)+\frac{3L\gamma^{2}}{2G}\sigma^{2}$
$+\gamma(2+3L\gamma)\delta_{1}+\gamma\zeta(N).$
@@
$E\left[\hat{f}\left(x^{T}\right)-\hat{f}\left(x^{0}\right)\right]$
$\leqslant-\frac{\gamma}{2}\left(\left(1-4\delta_{2}\right)-3L\gamma(1+2\delta_{2})\right)\sum\limits_{t=0}^{T-1}\mathbb{E}\left\|\nabla f\left(\overline{x}^{t}\right)\right\|^{2}$
$\quad+\frac{9\delta_{2}L\gamma^{2}(1+3L\gamma)(l-1)}{2}\sum\limits_{t=0}^{T-1}\mathbb{E}\left\|\nabla f(\overline{x}^{t})\right\|^{2}+\frac{9\delta_{1}L\gamma^{2}(1+3L\gamma)(l-1)T}{2}$
$\quad+3L^{2}\gamma^{3}(1+3L\gamma)\sigma^{2}(l-1)T+\frac{3L\gamma^{2}T}{2G}\sigma^{2}+\gamma(2+3L\gamma)\delta_{1}T+\gamma\zeta(N)T$
@@
$\quad+3L^{2}\gamma^{3}(1+3L\gamma)\sigma^{2}(l-1)T+\frac{3L\gamma^{2}T}{2G}\sigma^{2}+\frac{9\delta_{1}L\gamma^{2}(1+3L\gamma)(l-1)T}{2}$
$\quad+\gamma(2+3L\gamma)\delta_{1}T+\gamma\zeta(N)T.$
@@
$1-4\delta_{2}-3L\gamma(1+2\delta_{2})-9L\gamma\delta_{2}(1+3L\gamma)(l-1)$ $\geqslant$ $\frac{1}{2}-\frac{15}{4}L\gamma$
$-\frac{9}{8}L\gamma\left(1+\frac{3}{25(l-1)}\right)(l-1)$
@@
$+\frac{5}{2}\gamma\delta_{1}T+6L\gamma^{2}\delta_{1}(l-1)T+\gamma\zeta(N),$
@@
$+13\delta_{1}+30L\gamma(l-1)\delta_{1}+5\zeta(N),$
@@
In this remark we want to explain why $\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}\left\|\nabla f\left(\frac{1}{G}\sum\limits_{i\in\mathcal{G}}x_{i}^{t}\right)\right\|^{2}$, that we choose as a criterion in Theorem G.3 is consistent. When we consider iterations, when Algorithm 7 performs aggregation, we assign weights of honest devices, i.e. devices, which send honest stochastic gradient $g$ with $\mathbb{E}[g]=\nabla f(x)$ equal to $\frac{1}{G}$, and weights of other devices (who acts as Byzantine) is equal to 0. And this layout gives an estimate that is an upper bound of true $\omega$ realization (G.2). In other words, we say, that if algorithm at each aggregation round take average of points from devices, which all previous local round act like honest it would not be better than for convergence, than real iteration of the algorithm. In such a way, we a not interesting in the points of devices, who perform even one Byzantine-like iteration in the local round. Now it is clear why we have to weaken the assumption about at least one honest device at each iteration, not necessary the same: we request at least one honest device at each local round, not necessary the same in different rounds.
@@
$\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}\left\|\nabla f(\overline{x}^{t})\right\|^{2}=$ $\mathcal{O}\left(\frac{\mathbb{E}\left[\hat{f}\left(x^{0}\right)-\hat{f}\left(\hat{x}^{*}\right)\right]lL}{T}+\frac{\sigma\sqrt{\mathbb{E}\left[\hat{f}\left(x^{0}\right)-\hat{f}\left(\hat{x}^{*}\right)\right]L}}{\sqrt{TG}}+\delta_{1}+\zeta(N)\right).$
@@ Appendix H Partial participation
In the previous section, we discussed an important regime in distributed systems known as the local approach. Another widely used scenario in distributed learning is partial participation, which can be advantageous in various practical setups [yang2021achieving, wang2022unified, li2022federated]. In this section, we extend our AutoBant algorithm to support partial participation in federated learning. Below, we provide a formal description of the AutoBant with Partial Participation method (Algorithm 8).
@@
Under Assumptions 3.1, 3.2(c), 3.3, 3.4 with $\delta_{2}\leqslant 0.25$, 3.5, for solving the problem (1), after $T$ iteration of Algorithm 2 in partial participation scenario with $\gamma\leqslant\frac{1}{15L}$, it implies
$\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}\|\nabla f(x^{t})\|^{2}$ $\leqslant$ $\frac{4\mathbb{E}\left[\hat{f}\left(x^{0}\right)-\hat{f}\left(\hat{x}^{*}\right)\right]}{\gamma T}+3\delta_{1}+\frac{6L\gamma}{\widetilde{G}}\sigma^{2}+4\zeta(N),$
@@
$\hat{f}(x^{t+1})$ $=$ $\min_{\omega\in\Delta_{1}^{n}}\hat{f}\left[x^{t}-\gamma\sum_{i\in\mathcal{W}(t)}\omega_{i}g_{i}^{t}\right]$
$\leqslant$ $\hat{f}\left[x^{t}-\frac{\gamma}{G(t)}\sum_{i\in\mathcal{G}(t)}g_{i}^{t}\right]$
[…]
@@
$\mathbb{E}\hat{f}(x^{t+1})$ $\leqslant$ $\mathbb{E}\hat{f}(x^{t})-\left\langle\nabla\hat{f}(x^{t}),\frac{\gamma}{G(t)}\sum_{i\in\mathcal{G}(t)}\nabla f_{i}(x^{t})\right\rangle$
$+\frac{L\gamma^{2}}{2}\mathbb{E}\left\|\frac{1}{G(t)}\sum_{i\in\mathcal{G}(t)}g_{i}^{t}\right\|^{2}$
@@
$+\frac{3\gamma}{2}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}\right)+\frac{L\gamma^{2}}{2}\mathbb{E}\left\|\frac{1}{G(t)}\sum_{i\in\mathcal{G}(t)}g_{i}^{t}\right\|^{2}$
[…]
$+\frac{3L\gamma^{2}}{2}\left\|\frac{1}{G(t)}\sum_{i\in\mathcal{G}(t)}(\nabla f(x^{t})-\nabla f_{i}(x^{t}))\right\|^{2}$
$+\frac{3L\gamma^{2}}{2}\mathbb{E}\left\|\frac{1}{G(t)}\sum_{i\in\mathcal{G}(t)}(\nabla f_{i}(x^{t})-g_{i}^{t})\right\|^{2}$
$+\frac{3L\gamma^{2}}{2}\left\|\frac{1}{G(t)}\sum_{i\in\mathcal{G}(t)}\nabla f(x^{t})\right\|^{2}.$
@@
$+\frac{3L\gamma^{2}}{2}\left(\frac{1}{G(t)}\sum_{i\in\mathcal{G}(t)}\left\|\nabla f(x^{t})-f_{i}(x^{t})\right\|^{2}\right$.
$\left.+\frac{1}{(G(t))^{2}}\sum_{i\in\mathcal{G}(t)}\mathbb{E}\left\|\nabla f_{i}(x^{t})-g_{i}^{t}\right\|^{2}\right)+\frac{3L\gamma^{2}}{2}\left\|\nabla f(x^{t})\right\|^{2}$
[…]
$+\frac{3L\gamma^{2}}{2}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}+\frac{\sigma^{2}}{G(t)}\right)+\frac{3L\gamma^{2}}{2}\left\|\nabla f(x^{t})\right\|^{2}$
@@
$+\frac{2\gamma}{2}(1+3L\gamma)\delta_{1}+\frac{3L\gamma^{2}}{2G(t)}\sigma^{2}+\gamma\zeta(N).$
@@
$\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}\|\nabla f(x^{t})\|^{2}$ $\leqslant$ $\frac{4\mathbb{E}\left[\hat{f}\left(x^{0}\right)-\hat{f}\left(\hat{x}^{*}\right)\right]}{\gamma T}+3\delta_{1}+\frac{6L\gamma}{\widetilde{G}}\sigma^{2}+4\zeta(N).$
@@
$\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}\left\|\nabla f(x^{t})\right\|^{2}=$ $\mathcal{O}\left(\frac{\mathbb{E}\left[\hat{f}\left(x^{0}\right)-\hat{f}\left(\hat{x}^{*}\right)\right]L}{T}+\frac{\sigma\sqrt{\mathbb{E}\left[\hat{f}\left(x^{0}\right)-\hat{f}\left(\hat{x}^{*}\right)\right]L}}{\sqrt{T\widetilde{G}}}+\delta_{1}+\zeta(N)\right).$
@@ Appendix I SimBant
We also propose another method for dealing with Byzantine attacks - SimBant. Here we use the concept of the trial function in a different way: the key is that we look not at the loss function $\hat{f}$, as in Algorithms 1, 2, but at the output of the model $m$ on the trial data $\hat{D}$. In particular, the calculation of trust scores $w_{i}$ is now based on how the outputs of the model parameters obtained on the server (a guaranteed honest device, without losing generality, we can assume that it has index 1) and on the device are similar to each other. The server is validated in pairs with each device and the trust scores are calculated for each device relative to the prediction classes based on how similar they are between the server and the device. And only on the basis of the trust scores, we give a weights to devices’ trained models and calculate the new state of the general model. Let us denote by $g^{t}$ the model that the server sends to the devices at step $t$, $g_{i}^{t},i\in\{2,\ldots,n\}$ - devices’ trained models, $g_{1}^{t}$ the trained server model. Then the trust score for $i$-th device is function $)\alpha_{i}\rightarrow sim(m(x^{t}-\gamma g_{i}^{t},\mathcal{\hat{D}}),m(x^{t}-\gamma g_{1}^{t},\mathcal{\hat{D}}$)), which measures the similarity of predictions of server and device models, where $x^{t}-\gamma g_{i}^{t}$ and $x^{t}-\gamma g_{1}^{t}$ are new parameters on the $i$-th device and the server, respectively, simsim is responsible for the measure of the similarity of the outputs. We do not provide the theory for this method but validate it in practice (see Section 6).
@@
10:  $\omega_{i}^{t}=(1-\beta)\omega_{i}^{t-1}+\beta\frac{sim(m(x^{t}-\gamma g_{i}^{t},\hat{D}),m(x^{t}-\gamma g_{1}^{t},\hat{D}))}{\sum\nolimits_{j=1}^{n}sim(m(x^{t}-\gamma g_{j}^{t},\hat{D}),m(x^{t}-\gamma g_{1}^{t},\hat{D}))}$
11:  $x^{t+1}=x^{t}-\gamma\sum\nolimits_{i=1}^{n}\omega_{i}^{t}g_{i}^{t}$